# Learning What Matters: Prioritized Concept Learning via Relative Error-driven Sample Selection

### v1

@@ B.1 Ablation Studies
(c) Selection Gap $\gamma$*$BatchSize$
@@
We conduct further ablations studies in Fig. 9 on effect of hyperparameters. All experiments use LLaVA-v1.5-7B on the LLaVA-665K dataset with 20% sampling ratio and accuracy as the objective. Fig. 9(a) shows the effect of different warm-up ratios relative to the total training data pool size. Our results show that a 9% warm-up ratio achieves the best performance, as it strikes a balance between preparing the model adequately and leaving enough room for our iterative Prioritized concept learning strategy to select informative samples. The 20% warm-up ratio (i.e using only warmup selected samples to entirely train the model eleminating our Prioritized concept learning strategy completely), results in significantly reduced performance in overall relative score. Next, Fig. 9(b) shows the effect of varying the number of clusters $K$ used for our concept categorization module (Section 3.1). Using too few skill-clusters reduces skill diversity and leads to lower purity (in terms of skill types) within given cluster, while too many clusters result in redundant clusters of the same skill category and insufficient samples per cluster to yield credible accuracy estimates. We find that using approximately 1,000-2000 clusters strikes the best balance and yields optimal performance. Finally, Fig. 9(c) shows the influence of the selection gap i.e $\gamma$$BatchSize$ (see definition in Appendix 4). We find that the model is particularly sensitive to small gaps; for instance, a gap size of 1,500 leads to a rapid performance decrease. Smaller gaps cause the model to switch too soon, not allowing it to learn the selected concepts sufficiently.
How important is the order of skill acquisition?
@@
Unlike prior methods that focus solely on selecting which samples to use (Lee et al., 2024; Wu et al., 2025), PROGRESS also controls when to introduce them during training (Section 3.2)—effectively guiding both skill selection and the order of acquisition. To assess the importance of learning order, we ablate this component by randomly shuffling the data selected by PROGRESS and training the model without respecting the intended sequence. Even with the same data, training in a random order leads to a noticeable performance drop—from 98.8% to 94.6%—highlighting that when to introduce concepts is just as important as what to learn.
Figure 10: Wall-clock time comparison. The time cost is measured in hours of running time on a computing node with 4 × A100 GPUs.
@@
Wall-clock Time Comparison.
We present wall-clock time analysis with Qwen2-VL-7B on the LLaVA-1.5 dataset using 20% selected data for training. We compare average relative performance (Rel.) against wall-clock time for two of the highest-performing baselines—COINCIDE (Lee et al., 2024) and Random. PROGRESS (Accuracy as Objective) achieves relative performances of 96.3%, 97.1%, 98.2%, 99.1%, and 100% with wall-clock times of 1.5, 1.75, 2.5, 4, and 5.67 hours, respectively—making PROGRESS Pareto superior to COINCIDE. Full fine-tuning on entire dataset (with 100% data) takes about 9 hours. Note that the plot includes only actual data selection and model training time for both methods and excludes the time to train additional reference VLM model which COINCIDE needs for clustering and also accuracy estimation inference time for cluster accuracy evaluation using an LLM judge for our method. For our approach, accuracy estimation using LLM overhead is minimal when leveraging online concurrent APIs—and can be further reduced by using loss as the objective variant of our method which provides state-of-art performance and does not require cluster level accuracy estimation (see performance in Table 1 in main manuscript).
@@
B.2 Details for Analysis in main manuscript- How does the benchmark difficulty and data frequency impact performance?
@@
0:  Training dataset $\mathbb{U}=\{x_{1},\ldots,x_{N}\}$ where $x_{i}=(I_{i},Q_{i})$; set of benchmarks =$\mathcal{B}=\{B_{1},\ldots,B_{M}\}$ with $M$ different benchmarks.
@@
3:  $n_{j}\leftarrow 0,\ \forall j\in\{1,\ldots,M\}$Initialize match counts for each benchmark}
@@
B.2.1 Details for Benchmark Difficulty Analysis
@@
[…]
@@
B.2.2 Details for Data frequency Analysis
@@
frequency=# matched samplestotal training $\text{frequency}=\frac{\text{\# matched samples}}{\text{total training samples}}.$
@@
B.3 Details for Analysis - What Skills Does the Model Prioritize and When?
@@
0:  Training dataset $\mathbb{U}=\{x_{1},\ldots,x_{N}\}$ where $x_{i}=(I_{i},Q_{i})$; Clusters $\mathcal{C}=\{C_{1},\dots,C_{K}\}$; samples from MME benchmark =$\mathcal{B}=\{b_{1},\dots,b_{M}\}$ with ability labels; similarity threshold $\alpha=0.9$; top-$k$ nearest neighbors ($k=5$)
@@
Appendix C Analysis
C.1 Word Cloud Visualization of Skill Clusters
@@
C.2 Skill-level Diversity in Selected Sample Distribution
To better understand the selection behavior across methods, we follow protocol in previous work (Lee et al., 2024) and analyze the number of selected samples from each task in the Vision-Flan-191K dataset. Figure 12 shows the task-wise sample distribution for PROGRESS and several baseline approaches.
We observe that methods relying on single static scoring functions—such as CLIP-Score, EL2N, Perplexity, and Self-Sup—tend to exhibit strong sampling bias, disproportionately selecting from a small subset of tasks while neglecting others. This narrow focus often overlooks important data modes, leading to poor generalization—a limitation also noted in prior work (Lee et al., 2024).

### v2

@@ 7.2 Ablation Studies
(c) Selection Gap $\gamma*$BatchSizeBatchSize
@@
We conduct further ablations studies in Fig. 10 on effect of hyperparameters. All experiments use LLaVA-v1.5-7B on the LLaVA-665K dataset with 20% sampling ratio and accuracy as the objective. Fig. 10(a) shows the effect of different warm-up ratios relative to the total training data pool size. Our results show that a 9% warm-up ratio achieves the best performance, as it strikes a balance between preparing the model adequately and leaving enough room for our iterative Prioritized concept learning strategy to select informative samples. The 20% warm-up ratio (i.e using only warmup selected samples to entirely train the model eleminating our Prioritized concept learning strategy completely), results in significantly reduced performance in overall relative score. Next, Fig. 10(b) shows the effect of varying the number of clusters $K$ used for our concept categorization module (Section 3.1). Using too few skill-clusters reduces skill diversity and leads to lower purity (in terms of skill types) within given cluster, while too many clusters result in redundant clusters of the same skill category and insufficient samples per cluster to yield credible accuracy estimates. We find that using approximately 1,000-2000 clusters strikes the best balance and yields optimal performance. Finally, Fig. 10(c) shows the influence of the selection gap i.e $\gamma$ *BatchSizeBatchSize (see definition in Appendix 6.2). We find that the model is particularly sensitive to small gaps; for instance, a gap size of 1,500 leads to a rapid performance decrease. Smaller gaps cause the model to switch too soon, not allowing it to learn the selected concepts sufficiently.
Why PROGRESS Works: Representativeness of Selected Subsets and Order of Skills
Representativeness of the Selected Subset. We evaluate how well the subsets selected by PROGRESS and COINCIDE reflect the full dataset using PCA-based analyses. We first concatenate the DINO and BERT features and fit a single Principal Component Analysis (PCA) model on the entire dataset, reducing the dimensionality to 50 components. This global PCA transform is then applied consistently to the full dataset and to the samples selected by both methods. In this 50-dimensional PCA space, we compute two measures of representativeness:
  • Mean Nearest-Neighbor (NN) Distance. To estimate how densely each subset covers the data manifold, we build a FAISS IndexFlatL2 index containing all PCA-transformed points from the full dataset. For every point in a subset, we query this index to find its closest neighbor within the full dataset and measure that Euclidean distance. Averaging these distances yields a single value that reflects how close the subset stays to the dense, frequently occurring regions of the distribution. Lower values indicate that the subset lies in typical, well-populated regions rather than isolated or sparsely sampled areas. PROGRESS achieves a mean NN distance of ||0.083||, substantially lower than COINCIDE’s ||0.198||, indicating that PROGRESS selects samples closer to the core of the dataset.
    While mean NN distance is intuitive and meaningful, it can be minimized by degenerate subsets that concentrate in a few high-density regions and fail to preserve the global structure of the distribution. For this reason, we also report a complementary Variance Coverage metric.
  • Variance Coverage. PCA orders directions in the feature space by how much variation they capture: the leading components encode the dominant structural patterns of the dataset, while later components reflect finer-grained variations. To assess how well each subset preserves this structure, we compare its variance along every principal component to the corresponding variance in the full dataset. For each component, we compute $\text{ratio}=\frac{\text{variance (subset)}}{\text{variance (full)}},$ which measures how much of the dataset’s variation along that axis the subset retains. We summarize this comparison in two ways. First, we compute a Weighted Variance Retained score by weighting each variance ratio by the component’s explained variance in the full dataset and summing across all 50 components. This yields a single measure of how much of the dataset’s overall variability the subset preserves. PROGRESS retains more variance than COINCIDE (38.7% vs. 34.2%). Second, we directly examine the variance ratios for the top principal components, which capture the major structural axes of the dataset. High coverage on these components indicates that a subset spans the most important modes of variation rather than collapsing onto a narrower region. PROGRESS covers a substantially larger fraction of these leading components compared with COINCIDE (77–88% vs. 48–72%).
These results show that PROGRESS spans the overall data distribution more effectively, while COINCIDE selects a more concentrated and less representative subset.
@@
How important is the order of skill acquisition? Unlike prior methods that focus solely on selecting which samples to use [19, 45], PROGRESS also controls when to introduce them during training (Section 3.2)—effectively guiding both skill selection and the order of acquisition. To assess the importance of learning order, we ablate this component by randomly shuffling the data selected by PROGRESS and training the model without respecting the intended sequence. Even with the same data, training in a random order leads to a noticeable performance drop—from 98.8% to 94.6%—highlighting that when to introduce concepts is just as important as what to learn.
@@
Figure 11: Wall-clock time comparison on Vision-Flan dataset.
7.3 Wall-clock Time Comparison
We compare average relative performance (Rel.) against wall-clock time on the Vision-Flan dataset in Fig. 11. PROGRESS (Accuracy as Objective) achieves relative performances of 89.4%, 93.3%, 99.0%, 100.2% and 101.8% with wall-clock times of 0.69, 0.89, 1.21, 1.65 and 2.52 hours. Full fine-tuning on entire dataset takes about 2.56 hours. Remarkably, our method exceeds 100% relative performance (i.e surpasses vanilla full data fine-tuning) in just 1.65 hours—including both data selection and model training time—needing only 64% of the time required for full dataset fine-tuning.
Table 9: Time breakdown comparison: COINCIDE vs PROGRESS vs Full Data Training. We use official code to produce this time analysis
Stage Time Note Full Data Training 540 min (9h) Training on 100% data on 4×A100 COINCIDE Manual Inspection Unknown Manual MSA layer selection to extract features, needs ground-truth annotations Feature Extraction 280 mins TinyLLaVA feature extraction on 2×A100 Clustering 50 mins With faiss-gpu KNN on single A100. COINCIDE requires large memory to store 20,480 dim. features from multiple MSA layers. Training 150 mins Training on 20% data on 4×A100 Total 480 min (8h) PROGRESS Feature Extraction 30 mins Sentence-BERT (all-MiniLM-L6-v2) + DINO-v2-base on single A100 Clustering 30 mins With faiss-gpu KNN on single A100 Annotation Decision 130 mins Model self-evaluates and decides what to annotate next. 13 mins × 10 rounds Training 150 mins Training on 20% data on 4×A100 Total 340 min (5.67h)
7.4 Stage-Wise Time Breakdown Analysis
Table 9 presents a detailed time breakdown comparing our PROGRESS method with the closest competitor COINCIDE. We use same GPU compute for both COINCIDE and our method for fair comparison. COINCIDE requires over 8 hours of total computation time (plus additional unknown manual inspection time) and needs an additional pre-trained TinyLLaVA VLM to extract and store features from multiple MSA layers (4th, 8th, 12th, 16th, 20th). For each sample, they need 4096×5 = 20,480 dimensional features, which demand substantial memory resources and significantly increases clustering time. In contrast, our PROGRESS method completes in only 5.67 hours without requiring any additional VLM models. By using lightweight uni-modal self-supervised feature extractors—combining DINO-v2 (1,024 dimensions) and Sentence-BERT (384 dimensions)—we achieve efficient clustering with only 1,408-dimensional features per sample. This represents a 14.5× reduction in feature dimensionality compared to COINCIDE, while still achieving better relative performance w.r.t COINCIDE using only 20% of the training data.
7.5 Annotation Time Analysis: Significance of Annotation Cost as the Primary Bottleneck
In this section, we estimate the overall annotation cost and show that it is the dominant bottleneck in scaling VLM training to larger datasets—a cost that our method drastically reduces by 80% (as it requirs only 20% data for training).
The LLaVA-665K dataset has 0.67 million samples comprising Human-curated data (OKVQA, A-OKVQA, OCRVQA, TextCaps) and Synthetic QA which still need human-annotated bounding box & object names (from COCO etc, see details in [25, 23]) which are then provided to an LLM to generate Q/A.
Prior studies [8] estimate it typically takes 10.3 sec for a human (Mturk worker) to annotated 1 sample (which consists Image & 4-5 Q/A pairs). So, to annotate 0.67 million samples in LLaVA dataset it will take 1902 hours (0.67 M * 10.3 / 3600).
Estimated annotation time:
• Full-data training (100% of 0.67M samples): $\sim$1902 hr
• PROGRESS (20% of 0.67M samples): $\sim$380 hr
Overall time (annotation + training):
• Full-data finetuning: 1902 hr (annotation) + 9 hr (training)
• PROGRESS : 380 hr (annotation) + 5.7 hr (training)
Conclusion. Annotation is by far the dominant cost and major bottleneck in VLM training, especially when we scale to even larger datasets. By reducing annotation to only 20%, PROGRESS drastically cuts the primary bottleneck and offers a scalable path for training on even larger datasets.
7.6 Scaling data improves PROGRESS even more than vanilla Full-Finetuning
We point to results shown in Table 10 and Fig 7 where scaling data size to higher percentages 32%, 64%, our method outperforms full-finetuning (which uses 100% data) by larger and larger margin (see Rel. Score) as data % is scaled up.
Table 10: Scaling performance of PROGRESS .
Data Used Rel. Score (%) Full Finetune 100% 100 PROGRESS 4.2% 89.4 8.3% 93.3 16.7% 99.0 32.0% 100.2 64.5% 101.8
Reason for better scaling performance- Our method removes redundancy & focuses on most informative samples that the model should learn next - naturally shifts attention toward skills that show strong learning potential, instead of spending excessive effort on skills it already performs well on or are too hard to learn at this instant of time.
@@ 7.6 Scaling data improves PROGRESS even more than vanilla Full-Finetuning
7.7 Details for Analysis in main manuscript- How does the benchmark difficulty and data frequency impact performance?
@@
0: Training dataset $\mathbb{U}=\{x_{1},\ldots,x_{N}\}$ where $x_{i}=(I_{i},Q_{i})$; set of benchmarks $\mathcal{B}=\{B_{1},\ldots,B_{M}\}$ with $M$ different benchmarks.
@@
[…]
@@
7.7.1 Details for Benchmark Difficulty Analysis
@@
Prior work has shown that human intuition about task difficulty may not align with a model’s difficulty as defined in its feature or hypothesis space [38]. Therefore, we use the model’s own performance as a proxy for determining benchmark difficulty. Specifically, we use the performance of full-dataset fine-tuned LLaVA-v1.5-7B (i.e., Row 0 in Table 1) as reference to determine difficulty of benchmark—benchmarks with higher performance are considered easier. We define benchmark difficulty for a given benchmark as (100Performance of full fine-tuned LLaVA-1.5 on Benchmark)/100(100-$\text${Performance of full fine-tuned LLaVA-1.5 on Benchmark})/100. This gives us a difficulty measure for each benchmark normalized between $[0,1]$  [^3] For MME, where the full score is not out of 100, we normalize the score by dividing it by the maximum score (2800), the difficulty is computed as (1MME Score/2800)(1-$\text${MME Score}/2800).
@@
7.7.2 Details for Data frequency Analysis
@@
$\text{frequency}=\frac{\text{\# matched samples}}{\text{total training samples}}.$
@@
7.8 Details for Analysis - What Skills Does the Model Prioritize and When?
@@
0: Training dataset $\mathbb{U}=\{x_{1},\ldots,x_{N}\}$ where $x_{i}=(I_{i},Q_{i})$; Clusters $\mathcal{C}=\{C_{1},\dots,C_{K}\}$; samples from MME benchmark $\mathcal{B}=\{b_{1},\dots,b_{M}\}$ with ability labels; similarity threshold $\alpha=0.9$; top‑$k$ nearest neighbors ($k=5$)
@@
8 Analysis
8.1 Word Cloud Visualization of Skill Clusters
@@
8.2 Skill-level Diversity in Selected Sample Distribution
To better understand the selection behavior across methods, we follow protocol in previous work [19] and analyze the number of selected samples from each task in the Vision-Flan-191K dataset. Figure 13 shows the task-wise sample distribution for PROGRESS and several baseline approaches.
We observe that methods relying on single static scoring functions—such as CLIP-Score, EL2N, Perplexity, and Self-Sup—tend to exhibit strong sampling bias, disproportionately selecting from a small subset of tasks while neglecting others. This narrow focus often overlooks important data modes, leading to poor generalization—a limitation also noted in prior work [19].
@@ 8.2 Skill-level Diversity in Selected Sample Distribution
9 Limitations
While PROGRESS effectively orders and prioritizes more informative skills, it randomly samples within each selected skill cluster without ranking samples by usefulness. Additionally, the accuracy-based variant incurs extra inference time to compute skill-level progress (see Appendix 6), though our loss-based variant avoids this issue. However, overall, PROGRESS outperforms prior approaches while requiring no additional reference VLM and significantly less supervised data.
10 LLM Usage Disclosure
The LLM was primarily used for language refinement rather than any content generation—all experimental designs, results, analyses, and scientific contributions are original work by the authors. The LLM assistance was limited to editorial improvements such as fixing grammatical errors, suggesting clearer phrasing for complex technical concepts, and ensuring consistency in terminology throughout the manuscript. No experimental results, mathematical derivations, or scientific claims were generated by the LLM. All factual statements, citations, and technical content were independently verified by the authors.